Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
Answer by ptaylor(2048)

Let x=number of nickels
y=number of dimes
Lets deal in pennies:
We are told that:
5x+10y=565 divide each term by 5
x+2y=113----------------------------------eq1
and
5(x+8)+10*2y=1045=
5x+40+20y=1045 subtract 40 from each side
5x+20y=1005 divide each term by 5
x+4y=201-----------------------------------------eq2
subtract eq1 from eq2:
2y=88
y=44-------------------------- number of dimes
substitute into eq1:
x+88=113
x=25-------------------------number of nickels
CK
5*25+10*44=565
125+440=565
565=565
and
5*(25+8)+10*88=1045
165+880=1045
1045=1045

Hope this helps----ptaylor